Grade 12th passGeneral Physics

A uniform shell of radius 4r/3 and mass M has a point mass M at its centre . Two points A and B are at distance r and 2r from the centre inside and outside the shell resp .Choose the correct option
a ) the ratio of the gravitational potential at A to that at B is 7;:4
b ) the gravitational interactional energy between point mass and shell is – 3GM2 /4r
c ) escape velocity under root 3GM/R

Chanchal
10 Years agoGrade 12th pass

1 Answer

Vikas TU
10 Years ago
  1. At A position,
Gravitational potential = GM/r + 3GM/4r = 7GM/4r
At B position,
Gravitational potential = GM/2r + GM/2r = GM/r
Ratio => 7/4 = > 7:4